# Equivariant Symmetry-Aware Head Pose Estimation for Fetal MRI

### v4

@@ Appendix F Baseline methods
EquiTrack [billot2024]
This method first uses a “denoiser” CNN to remove intensity differences between input and template volumes, followed by a SE(3)-equivariant network that learns matching landmarks. Here, we re-implement EquiTrack to guarantee fair comparison with our method. First, we retrain the denoiser network to remove spin history artifacts in addition to standard intensity transformations in brain MRI volumes, such as bias field and noise. Second, we use RANSAC [fischler1981] to compute the optimal rigid transform to further boost robustness to disruptive intensity perturbations (e.g., spin history artifacts).
Fetal-Align [hoffmann2021]
The original implementation uses standard image processing techniques (i.e., detection of maximally stable extremal regions [matas2004]) conditioned on GA, rather than deep learning, to estimate brain and eye masks in fetal brain volumes. We find that this approach to segmentation becomes unstable when applied to clinical volumes. For fair evaluation of landmark-based pose estimation on our data, we use the brain and eye masks predicted by our segmentation U-Net trained on simulated spin history artifacts. When eye detection fails (a common occurrence when the artifact obstructs the eyes in navigator volumes), we return the identity transformation as an estimate of rotation.
3DPose-Net [salehi2019]
We train this network with the $\mathcal{L}_{1}$ loss function instead of the geodesic loss proposed in the original work. We find that this training objective provides better performance on our datasets. For our training experiments on Research-Fetal, we use the same augmentations as that of E(3)-Pose. For our training experiments on dHCP, we tune the data augmentation parameters such that $\log{\gamma}$ is sampled uniformly from [-3,-1] and voxel size is sampled uniformly from [3,9]mm.
6DRep [faghihpirayesh2023]
Similar to 3DPose-Net, we replace the geodesic loss function with the $\mathcal{L}_{1}$ loss function, which operates on the output and GT orthonormal basis directions. Again, we find that this modification improves the test-time performance on our data. We use the same training augmentations as 3DPose-Net.
RbR [gopinath2024]
We find that this method provides more stable and accurate pose estimates on fetal brain volumes that are preprocessed with the same steps as our E(3)-CNN, i.e., inputs that are center-of-mass aligned and scaled based on the size of the brain. Since center-of-mass alignment of input volumes eliminates the need for translation estimation, we keep the translation estimation of E(3)-Pose and predict only rotation with the output of RbR. We use the same training augmentations as 3DPose-Net. We use Adam optimization with learning rate $10^{-4}$ to train the network.
Appendix G Additional results
@@
We separately evaluate the accuracy of translation estimation with our segmentation network by providing the translation error $\hat{t}-t||_{2}$. We compare E(3)-Pose to baseline methods that estimate translation differently from E(3)-Pose, i.e., FireANTs and EquiTrack. Table 3 shows that E(3)-Pose outperforms both FireANTs and EquiTrack on navigator volumes. The presence of spin history artifacts breaks the assumption of similarity between the input and template volume intensity distributions. Template-free pose regression improves robustness to highly disruptive artifacts compared to the template-based alignment strategy.
@@
In this section, we provide additional example alignments produced by E(3)-Pose aand the baseline methods. Fig. 9 displays the same volumes aligned by the baseline methods that were omitted from Fig. 3 in the main paper, i.e., EquiTrack, 3DPose-Net, and 6DRep. Fig. 10 displays the same set of volumes aligned by all methods trained on dHCP. Fig. 11 replicates Fig. 3 for a different set of test volumes in each dataset.
Appendix H Sensitivity study
@@
Appendix I Ablation study
@@
Appendix J Simulation study
@@
Recall from Appendix A that every slice is prescribed according to $\tilde{P}_{k}\!=\!\hat{T}_{k}P_{k}$, where $\hat{T}_{k}$ is the estimated head pose in the preceding navigator volume. We compare $\hat{T}_{k}$ returned by E(3)-Pose to that of “motion-blind” slice prescription, where we set $\hat{T}_{k}\!=\!T_{k=0}$ for all $k\!\geq\!0$. To simulate inter-slice fetal brain motion for k>0k\!>\!0, we sample rigid motion $T_{k+1}T_{k}^{-1}$ from real fetal head motion trajectories, interpolated at time intervals of TR=3s to match the time between navigator volume acquisitions in the clinical setting [shoemake1985]. Our motion trajectory dataset consists of full time-series from subjects in Clinical-Young, since increased uterine mobility in younger fetuses produces wider ranges of motion for our analysis. We follow prior work [xu2022] in annotating motion trajectories in this dataset for the purposes of simulation. Specifically, we apply an existing fetal landmark detection algorithm to detect the fetal eyes and shoulders in every frame [diaz2025], manually correct algorithmically estimated landmarks, and finally compute the rigid motion between consecutive frames using detected landmarks.
We define the slice coverage distribution $p_{C}$ over the underlying brain volume on a 1mm resolution grid as the sum of the Gaussian kernels representing the point spread function [rousseau2006, kuklisovamurgasova2012, xu2022] of all slices in the stack. We then compute the coverage gap as the proportion of the brain volume where $p_{C}\!=\!0$.
@@
Automated prescription of diagnostic 2D slices in fetal brain MRI aims to improve spatial coverage of the brain and the radiological interpretability of the acquired stack of slices. In the main paper, we partially quantify these qualities with the coverage gap and slice obliqueness, respectively. In this section, we include two additional metrics that further assess the diagnostic potential of the simulated slice stacks. First, because uneven spatial coverage is correlated with poor coverage quality and coverage gaps, we calculate the coverage irregularity $\text{KL}(p_{C}||p_{U})$, where KL is the Kullback-Leibler divergence and $p_{U}$ is the uniform distribution over the underlying brain volume. Second, since slice obliqueness only quantifies the rotational incoherence of slices within each stack, we also assess translational incoherence by computing the slice offset, i.e., the distance in mm between the center voxels of the GT and prescribed slices. Fig. 14 demonstrates that in our simulations, slice prescription with E(3)-Pose significantly improves both additional quantitative metrics of diagnostic potential, compared to motion-blind prescription.
Appendix K Rotational object symmetries
@@
where $D^{l}:\text{SO(3)}\rightarrow\mathbb{R}^{2l+1\times 2l+1}$ is the Wigner-D matrix function of order $l$, and $D^{l}_{:,m}$ denotes its columns, indexed from $l\!\leq\!m\!\leq\!l$. Columns of the Wigner-D matrices are irreducible tensors with irreducible Wigner-D matrix representations [weiler2018, geiger2022]. Since $h(R)$ is parametrized in terms of irreducible tensors, it maintains equivariance under E(3). Therefore, the E(3)-CNN output is formulated as 1 vector and 2 irreducible tensors of order l=Nl\!=\!N and parity $p\!=\!\text{parity}(N)$.
@@
where we have used the property $Y_{m}(\theta^{\prime},\phi^{\prime})\!=\!\sum\limits_{-l\leq m\leq l}D^{l}_{m^{\prime}m}(R)Y^{l}_{m^{\prime}}(\theta,\phi)$.
@@
where we use the property $Y_{l}^{m}(\theta_{1}\!+\!\theta_{2},\phi)\!=\!e^{im\theta_{2}}Y_{l}^{m}(\theta_{1},\phi)$.
@@
Since Wigner-D matrices are a Fourier basis for functions defined on SO(3), the E(3)-CNN output represents a likelihood function $P$ over SO(3), where $P(R^{\prime})$ is monotonically related to the inner product between $h(R^{\prime})$ and $\hat{e}_{x}\oplus\hat{e}_{y}\oplus\hat{e}_{z}] [$lee2024]. It is possible to sample from $P$ by querying values on a predefined equivolumetric grid over SO(3) [gorski2005]. Indeed, if the E(3)-CNN output $\hat{e}_{x}\oplus\hat{e}_{y}\oplus\hat{e}_{z}]$ is equal to $h(R)$, then $P$ is nonzero and uniformly distributed over the set of poses that are symmetrically equivalent to $R$, and zero elsewhere.

### v5

@@ Appendix H Baseline methods
EquiTrack [3]
This method first uses a “denoiser” CNN to remove intensity differences between input and template volumes, followed by a SE(3)-equivariant network that learns matching landmarks for rigid registration. Here, we slightly modify EquiTrack to guarantee fair comparison with our method on Navigators. First, to account for spin history artifacts, we train the denoiser network using our data augmentation strategy, which simulates these artifacts. Second, to further boost robustness to disruptive intensity perturbations (e.g., spin history artifacts), we use RANSAC [17] to compute the optimal rigid transform from matching landmarks. We train the denoiser and equivariant network for 550 and 2500 epochs with batch sizes of 4 and 1, respectively. Although we observed no statistically significant differences in performance statistics in Research-Fetal and dHCP, our modifications led to significant improvements of 5.1 and 10.2 in rotation error in Clinical-Young and Navigators, respectively.
Fetal-Align [33]
The original implementation uses standard image processing techniques (i.e., detection of maximally stable extremal regions [49]) conditioned on GA, rather than deep learning, to estimate brain and eye masks in fetal brain volumes. We find that this approach to segmentation becomes unstable when applied to our data. For fair evaluation of landmark-based pose estimation, we use the brain and eye masks predicted by our segmentation U-Net trained on simulated spin history artifacts. When eye detection fails (a common occurrence when the artifact obstructs the eyes in navigator volumes), we return the identity transformation as an estimate of rotation. Our substitution of hand-crafted image features for deep learning-based segmentation produced improvements in rotation error of 68.9 and 22.4 in Research-Fetal and dHCP, respectively; the improvements in clinical volumes were larger.
3DPose-Net [60]
For our training experiments on Research-Fetal, we use the same augmentations as that of E(3)-Pose. For our training experiments on dHCP, we tune the data augmentation parameters such that $\log{\gamma}$ is sampled uniformly from [-3,-1] and voxel size is sampled uniformly from [3,9]mm. We train the network for 8000 epochs.
6DRep [16]
We use the same training augmentations as 3DPose-Net and train the network for 7500 epochs.
RbR [25]
We find that this method provides more stable and accurate pose estimates on fetal brain volumes that are preprocessed with the same steps as our E(3)-CNN, i.e., inputs that are center-of-mass aligned and scaled based on the size of the brain. Since center-of-mass alignment of input volumes eliminates the need for translation estimation, we keep the translation estimation of E(3)-Pose and predict only rotation with the output of RbR. We use the same training augmentations as 3DPose-Net. We use Adam optimization with learning rate $10^{-4}$ to train the network for 4000 epochs.
Figure 13: Example results for ablations trained on Research-Fetal. See Fig. 3 in the main paper for descriptions.
Appendix I Additional results
@@
We separately evaluate the accuracy of translation estimation with our segmentation network by providing the translation error $\hat{t}-t||_{2}$. We compare E(3)-Pose to baseline methods that estimate translation differently from E(3)-Pose, i.e., FireANTs and EquiTrack. Tables 3 and 6 show that E(3)-Pose outperforms both FireANTs and EquiTrack in Navigators. The presence of spin history artifacts breaks the assumption of similarity of the intensity distributions between the input and template volumes. Compared to the template-based alignment strategy, template-free pose regression improves robustness to highly disruptive artifacts.
Evaluation of segmentation
We quantify the performance of our segmentation network by measuring the overlap between predicted and ground-truth brain masks with the Dice-Sorensen Coefficient (DSC) [83]. Our segmentation network yields mean DSC of 0.96, 0.98, 0.94, and 0.89 on Research-Fetal, dHCP, Clinical-Young, and Navigators, respectively (Fig. 9, left). While brain segmentation is robust across datasets, larger errors tend to occur in Navigators, which pose a challenging domain gap from the training data. We further observe that rotation estimation using our E(3)-CNN remains robust to segmentation errors, particularly larger errors in Navigators (Fig. 9, top right). This result highlights the advantages of a neural architecture constructed with strong inductive biases, coupled with a cropping margin large enough for the input volume to capture relevant anatomy regardless of segmentation errors. However, translation estimation is sensitive to segmentation error, since it is directly computed from the predicted mask (Fig. 9, bottom right).
Subject-level analysis in Navigators
To provide further insight, we aggregate performance statistics for each subject in Navigators. Tables 4-6 show that E(3)-Pose consistently outperforms the baselines across subjects and training datasets. Tables 7 and 8 show that regardless of the training data, E(3)-Pose ranks first, compared to all ablations, in the greatest number of subjects. Importantly, we highlight that for several subjects (S2 and S6), E(3)-Pose significantly outperforms all ablations. From visual inspection, we observe that these cases exhibit higher inherent pose ambiguity.
@@
In this section, we provide additional example alignments. Fig. 10 displays the same volumes aligned by the baseline methods that were omitted from Fig. 3 in the main paper, i.e., EquiTrack, 3DPose-Net, and 6DRep. Fig. 11 displays the same set of volumes aligned by all methods trained on dHCP. Fig. 12 replicates Fig. 3 for a different set of test volumes in each dataset. Fig. 13 provides example alignments for E(3)-Pose and all ablations.
Appendix J Sensitivity study
@@
Appendix K Ablation study
Alternative loss functions
In this section, we explain why the selected loss functions for our ablation experiments are not invariant under $G_{\text{symm}}$. Regarding the loss function where we replace the term $|\sin{(\theta_{x})}|$ with $|\sin{(\theta_{x}/2)}|$, Appendix D shows that $|\sin{(\theta_{x}/2)}|$ is monotonically related to $\hat{e}_{x}\cdot e_{x}$, which may change in value under the transformation $e_{x}\!\rightarrow\!-e_{x}$, i.e., left-right reflection. Regarding the geodesic loss, which measures the geodesic distance $d_{\text{geo}}(\hat{R},R)$ between the predicted and ground-truth rotations $\hat{R},R\!\in\!\text{SO(3)}$, it requires converting the network outputs $(\hat{e}_{x},\hat{e}_{y},\hat{e}_{z})$ to a proper rotation $\hat{R}$ via SVD. Although this process is approximately invariant under left-right reflections when $(\hat{e}_{x},\hat{e}_{y},\hat{e}_{z})\!\approx\!(e_{x},e_{y},e_{z})$, this condition is not guaranteed, especially at the start of training.
Results on research data
@@
Appendix L Simulation study
@@
Recall from Appendix A that every slice is prescribed according to $\tilde{P}_{k}\!=\!\hat{T}_{k}P_{k}$, where $\hat{T}_{k}$ is the estimated head pose in the preceding navigator volume. We compare $\hat{T}_{k}$ returned by E(3)-Pose to that of “motion-blind” slice prescription, where we set $\hat{T}_{k}\!=\!T_{k=0}$ for all $k\!\geq\!0$. To simulate inter-slice fetal brain motion for k>0k\!>\!0, we sample rigid motion $T_{k+1}T_{k}^{-1}$ from real fetal head motion trajectories, interpolated at time intervals of TR=3s to match the time between navigator volume acquisitions in the clinical setting [63]. Our motion trajectory dataset consists of full time-series from subjects in Clinical-Young, since increased uterine mobility in younger fetuses produces wider ranges of motion for our analysis. We follow prior work [77] in annotating motion trajectories in this dataset for the purposes of simulation. Specifically, we apply an existing fetal landmark detection algorithm to detect the fetal eyes and shoulders in every frame [15], manually correct algorithmically estimated landmarks, and finally compute the rigid motion between consecutive frames using detected landmarks.
We define the slice coverage distribution $p_{C}$ over the underlying brain volume on a 1mm resolution grid as the sum of the Gaussian kernels representing the point spread function [59, 40, 77] of all slices in the stack. We then compute the coverage gap as the proportion of the brain volume where $p_{C}\!=\!0$.
@@
The primary aims of automated slice prescription are improvements in the spatial coverage of the brain and the radiological interpretability of the acquired stack of slices. In the main paper, we partially quantify these qualities with the coverage gap and slice obliqueness, respectively. In this section, we include two additional metrics that further assess the diagnostic potential of the simulated slice stacks. First, because uneven spatial coverage is correlated with poor coverage quality and coverage gaps, we calculate the coverage irregularity $\text{KL}(p_{C}||p_{U})$, where KL is the Kullback-Leibler divergence and $p_{U}$ is the uniform distribution over the underlying brain volume. Second, since slice obliqueness only quantifies the rotational incoherence of slices within each stack, we also assess translational incoherence by computing the slice offset, i.e., the distance in mm between the center voxels of the GT and prescribed slices. Fig. 16 demonstrates that in our simulations, slice prescription with E(3)-Pose significantly improves both additional quantitative metrics of diagnostic potential, compared to motion-blind prescription.
Appendix M Additional limitations and future work
Navigators dataset
Although we evaluate E(3)-Pose on data representative of automated slice prescription with respect to acquisition parameters and image quality, our analysis is limited to 9 pregnant participants who were recruited at the same institution and imaged on the same scanner. Furthermore, Navigators does not include sufficient representation of second-trimester fetuses. To investigate the broader applicability of our method for clinical translation, future work will extend our analysis on navigator volumes acquired in a larger, more diverse cohort of pregnant volunteers.
Symmetries
Beyond fetal imaging, it is interesting to investigate applications of our method under more relaxed reflectional and/or rigid symmetries [69], such as cardiac, lung, and adult brain imaging. Lastly, we expect that our method could be adapted to objects with more complex (e.g., $n$-fold rotations) symmetries (Appendix N).
Appendix N Rotational object symmetries
@@
where $D^{l}:\text{SO(3)}\rightarrow\mathbb{R}^{2l+1\times 2l+1}$ is the Wigner-D matrix function of order $l$, and $D^{l}_{:,m}$ denotes its columns, indexed from $l\!\leq\!m\!\leq\!l$. Columns of the Wigner-D matrices are irreducible tensors with irreducible Wigner-D matrix representations [71, 21]. Since $h(R)$ is parametrized in terms of irreducible tensors, it maintains equivariance under E(3). Therefore, the E(3)-CNN output is formulated as 1 vector and 2 irreducible tensors of order l=Nl\!=\!N and parity $p\!=\!\text{parity}(N)$.
@@
where we have used the property $Y_{m}(\theta^{\prime},\phi^{\prime})\!=\!\sum\limits_{-l\leq m^{\prime}\leq l}D^{l}_{m^{\prime}m}(R)Y^{l}_{m^{\prime}}(\theta,\phi)$.
@@
where we use the properties $D^{l}(R_{1})D^{l}(R_{2})\!=\!D^{l}(R_{1}R_{2})$ for any $R_{1},R_{2}\!\in\!$ SO(3) and $Y_{l}^{m}(\theta_{1}\!+\!\theta_{2},\phi)\!=\!e^{im\theta_{2}}Y_{l}^{m}(\theta_{1},\phi)$.
@@
Since Wigner-D matrices are a Fourier basis for functions defined on SO(3), the E(3)-CNN output represents a likelihood function $P$ over SO(3), where $P(R^{\prime})$ is monotonically related to the inner product between $h(R^{\prime})$ and $\hat{e}_{x}\oplus\hat{e}_{y}\oplus\hat{e}_{z}] [$43]. It is possible to sample from $P$ by querying values on a predefined equivolumetric grid over SO(3) [26]. Indeed, if the E(3)-CNN output $\hat{e}_{x}\oplus\hat{e}_{y}\oplus\hat{e}_{z}]$ is equal to $h(R)$, then $P$ is nonzero and uniformly distributed over the set of poses that are symmetrically equivalent to $R$, and zero elsewhere.